solve assume a linear relationship fits each set of data
a long distance telephone comany advertised the following rates:
5 minutes (m) for just 0.85c (p) and 10 minutes for just $1.10 use the ordered pairs (m,p)
1.find the linear equation for these data points
2.use this linear equation to find the cost of a 20-minute phone call.

Answer by stanbon(75887)

solve assume a linear relationship fits each set of data
a long distance telephone comany advertised the following rates:
5 minutes (m) for just 0.85c (p) and 10 minutes for just $1.10 use the ordered pairs (m,p)
--------------
You have two points: (5,85) and (10,110)
slope = (110-85)/(10-5) = 25/5 = 5
intercept = b
85 = 5*5 + b
b = 60
------------------------

1.find the linear equation for these data points
p(m) = 5m + 60
------------------------
2.use this linear equation to find the cost of a 20-minute phone call.
p(20) = 5*20 + 60
p(20) = $1.60
=======================
Cheers,
Stan H.